# Yue Wanga and Jonas Berxa‡

### v1

@@ Conclusions
As a next step, it would be interesting to extend this framework to more complex geometries (e.g., non-circular confinements, porous media), incorporate hydrodynamic interactions, or study mixtures of active and passive tracers. Experimental validation using colloidal rollers or bacterial baths could further confirm the predictive power of braid measures. Beyond microfluidic mixing, the braiding framework established here may inform the design of reconfigurable active devices and our understanding of transport in natural and synthetic active systems, for instance through the determination of Lagrangian coherent structures acting as transport barriers 46, 47.
@@ Acknowledgements
The authors thank K. Proesmans for enlightening discussions.

### v2

@@ 6 Conclusions
As a next step, it would be interesting to extend this framework to more complex geometries (e.g., non-circular confinements, porous media), incorporate hydrodynamic interactions, or study mixtures of active and passive tracers. Experimental validation using colloidal rollers or bacterial baths could further confirm the predictive power of braid measures. We provided such a design for an experimental realisation of our system—based on artificial rollers—which can serve as a roadmap for pinpointing the proposed optimal mixing pattern. Beyond microfluidic mixing, the braiding framework established here may inform the design of reconfigurable active devices and our understanding of transport in natural and synthetic active systems, for instance through the determination of Lagrangian coherent structures acting as transport barriers 53, 54.
@@ Acknowledgements
The authors thank K. Proesmans for enlightening discussions. J.B. is supported by the Novo Nordisk Foundation with grant No. NNF18SA0035142.
@@
  • Bricard et al. 2015 A. Bricard, J.-B. Caussin, D. Das, C. Savoie, V. Chikkadi, K. Shitara, O. Chepizhko, F. Peruani, D. Saintillan and D. Bartolo, Nat. Commun., 2015, 6, 7470.
  • Bricard et al. 2013 A. Bricard, J.-B. Caussin, N. Desreumaux, O. Dauchot and D. Bartolo, Nature, 2013, 503, 95–98.
  • Zhang et al. 2020 B. Zhang, B. Hilton, C. Short, A. Souslov and A. Snezhko, Phys. Rev. Res., 2020, 2, 043225.
  • Wioland et al. 2016 H. Wioland, E. Lushi and R. E. Goldstein, New J. Phys., 2016, 18, 075002.
  • Zhang et al. 2020 B. Zhang, A. Sokolov and A. Snezhko, Nat. Commun., 2020, 11, 4401.
  • Hanke et al. 2013 T. Hanke, C. A. Weber and E. Frey, Phys. Rev. E, 2013, 88, 052309.
  • Imamura et al. 2023 S. Imamura, K. Sawaki, J. J. Molina, M. S. Turner and R. Yamamoto, Adv. Theory Simul., 2023, 6, 2200683.